Which of the following values represent the definite integral


"\intop\begin{matrix}
 \pi/4 \\
 0 
\end{matrix}" "e" tanθ sec2 θ dθ


1) e2


2) e - 1


3) e/2


4) e



1
Expert's answer
2021-08-23T16:24:26-0400

The correct answer is

2) e - 1

Solution;

Apply integration by substitution as follows;

Let "u=tan(\theta) \implies du=sec^2(\theta)d\theta"

This gives a ne lower bound and upper bound as follows;

Lower bound "u=tan(0) =0"

Upper bound "u=tan({\pi\over 4}) =1"

"\implies \int_0^{\pi\over 4}e^{tan(\theta)}sec^2(\theta)d\theta=\int_0^1e^udu=[e^u]_0^1=e^1-e^0=e-1"